Please help me with this question
A tank of 3500 fish have a mean of 1.2kg with a standard deviation of 0.225 kg.
Determine the number of these that will be:
A) More than 1.2kg
B) Less than 0.75kg
Regards
coolboy431

Answer by stanbon(75887)

A tank of 3500 fish have a mean of 1.2kg with a standard deviation of 0.225 kg.
Determine the number of these that will be:
A) More than 1.2kg
Find the z-value of 1.2
z(1.2) = (1.2-1.2)/0.225 = 0
P(x > 1.2) = P(z > 0) = 0.5000
------------------------------------
B) Less than 0.75kg
Find the z-value of 0.75
z(0.75) = (0.75-1.2)/0.225 = -0.45/0.225 = -2
P(x < 0.75) = P(z<-2) = 0.02275...
Cheers,
Stan H.